OpenStudy (anonymous):

Last one!. Simplify this expression. 2x3(5x3 - 7) A. 10x6 - 7 B. 10x6 - 14x3 C. 10x9 - 7 D. 10x9 - 14x3 Do i need to distribute it?

Parth (parthkohli):

Yes, you need to distribute.

OpenStudy (anonymous):

okay thanks!

Parth (parthkohli):

What do you get as the answer?

OpenStudy (anonymous):

10x6 - 14x3 ? i got confuse!.

OpenStudy (anonymous):

\[2x^3(5x^3 -7 )\\\\= 2*5x^{3+3}-2*7x^3\\\\= 10x^6 - 14x^3 \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

kk